OpenStudy (danieke):

medal and fan ? Jim runs 6 miles in 55 minutes. At the same rate, how many miles would he run in 44 minutes?

OpenStudy (anonymous):

\[\frac{6}{55}=\frac{x}{44}\] solve for \(x\) via \[x=\frac{6\times 44}{55}\] cancel first, multiply last

OpenStudy (danieke):

4.8?

OpenStudy (anonymous):

\[\frac{6\times 4}{5}\] is a first step so yeah what you said

OpenStudy (danieke):

i got it thank you

OpenStudy (anonymous):

yw